# Efficient Data Collection for Robotic Manipulation via Compositional Generalization

### v1

@@ B-A Robot Platform
B-B Evaluation Protocol
@@
We vary the factor table height by adjusting the height of the mobile table the robot and third-person cameras are mounted on. In other words, this factor is varied in practice by varying the height of the robot’s table, not the object table (the surface that objects are being manipulated on). However, this still changes the relative height of the robot with respect to the object table, the same as if the height of the object table was adjusted instead. In Section V-C, Higher Table refers to the object table being higher relative to the robot, which was achieved by lowering the robot table.
@@
We collect demonstrations using a Meta Quest 2 VR headset for teleoperation. All demonstrations are collected by a single experienced human teleoperator for consistency. We collect our Stair dataset by starting at $f^{*}$, and then varying factors cyclically in the order object position, object type, container type, table height, table texture. The order in which we vary values for each factor is the same as in Fig. 7, from top to bottom.
B-C Training
@@
We use the same ResNet-34 diffusion goal-conditioned policy architecture used in the original BridgeData V2 experiments, except we condition on a history of 2 128x128 RGB image observations from a third-person view (in addition to a goal image from the same view), and use action chunking to predict the next 4 actions. However, during inference we noticed better performance by only executing the first predicted action, so we do this for all experiments unless otherwise stated. We share the same visual encoder across image observations in the history. We use the same data augmentation procedure used in the original BridgeData V2 experiments, which consists of random crops, random resizing, and color jitter.
@@
For policies using BridgeData V2 as prior data, we first pre-train a model on prior data using the original training hyperparameters for 2M gradient steps, and choose the checkpoint with the lowest validation action prediction mean-squared error as the initialization for later fine-tuning. During fine-tuning, we use the same pre-training hyperparameters, except we train for only 300K gradient steps and use the final checkpoint. When co-fine-tuning, we train on a mixture of 75% in-domain data and 25% prior data.
B-D R3M

### v2

@@ Appendix C Real Robot Experiments
C-A Robot Platform
We use Logitech C920 webcams for our main and secondary third-person cameras. Our setup also includes the wrist camera used in BridgeData V2, and we experimented with using it in addition to our main third-person camera. While we found that it improved overall robustness when training from scratch, particularly for some factors that wrist cameras provide invariance to (e.g., object position, table height), policies still benefited from data coverage for these factors. More importantly, we also found that it was incompatible with using BridgeData V2 as prior data, reducing performance when using prior data compared to not using prior data at all. We believe this is because only a small fraction of BridgeData V2 contains wrist camera data, as similar negative results have been observed in other work that use prior robotic datasets with a small proportion of wrist camera data [35]. As we were only able to achieve successful transfer in our experiments in Section V-C by using prior data without a wrist camera, we decided to omit the wrist camera in all our experiments.
C-B Evaluation Protocol
@@
Our factor table height refers to the height of the object table (where objects are manipulated on) relative to the robot. However, we vary this factor in practice by adjusting the height of the mobile table the robot and third-person cameras are mounted on. This still changes the relative height of the robot with respect to the object table, the same as if the object table changed height. For example, in Section V-C, Higher Table refers to the object table being higher relative to the robot, which was achieved by lowering the robot’s table.
@@
We collect demonstrations using a Meta Quest 2 VR headset for teleoperation. All demonstrations are collected by a single experienced human teleoperator for consistency. We collect our Stair dataset by starting at $f^{*}$, and then varying factors cyclically in the order object position, object type, container type, table height, table texture. The order we vary values for each factor is the same as in Fig. 7, from top to bottom.
C-C Training
@@
We use the same ResNet-34 diffusion goal-conditioned policy architecture from the original BridgeData V2 experiments, except we condition on a history of 2 128x128 RGB image observations from a third-person view (in addition to a goal image from the same view), and use action chunking to predict the next 4 actions. However, we noticed better performance during inference by only executing the first predicted action, so we do this for all experiments unless otherwise stated. We share the same visual encoder across image observations in the history. We use the same 7D action space as the original experiments (6D end-effector pose deltas, and open/close gripper). We use the same data augmentation from the original experiments, which consists of random crops, random resizing, and color jitter.
@@
For policies using BridgeData V2 as prior data, we first pre-train a model on BridgeData V2 using the original training hyperparameters for 2M gradient steps. As done in the original BridgeData V2 experiments, 10% of trajectories in this data is reserved for validation. We checkpoint every 50K steps, and choose the checkpoint with the lowest validation action prediction mean-squared error as the initialization for later fine-tuning. During fine-tuning, we use the same hyperparameters as during pre-training, except we train for only 300K gradient steps. Also, we do not use a validation set, and instead simply evaluate the final checkpoint. When co-fine-tuning, we train on a mixture of 75% in-domain data and 25% prior data.
Data Strategy L Train Method Bridge From Scratch Factor 1 Factor 2 Object Pos Object Type Container Type Table Height Table Tex Object Pos Object Type Container Type Table Height Table Tex Object/Container Pos N/A 7/9 6/9 2/9 6/9 N/A 5/9 2/9 0/9 3/9 Object Orientation 2/9 5/9 5/9 1/9 3/9 0/9 2/9 3/9 0/9 1/9 Overall 36/81 16/81
TABLE VI: Additional real robot pairwise composition results for our “put fork in container” task, with additional factors object/container position and object orientation. Similarly as in our main pairwise composition results in Table I, leveraging BridgeData V2 as prior data significantly improves composition for these factors compared to training from scratch.
Figure 16: Visualization of our additional real robot factors in BaseKitch. The top row shows our base factor values $f^{*}$. The other rows show all deviations from $f^{*}$ by one factor value.
C-D Pairwise Composition for Additional Factors
We conduct experiments for two new factors: object/container position and object orientation. We visualize these factors and their values in Fig. 16. Note that object/container position involves new positions of both the fork and container, with more significant changes compared to our previous values for object position. We extended the L dataset for our original factors with 10 demonstrations for each new factor value, re-trained policies on this extended dataset, and evaluated pairwise composition with these new factors. Unlike our previous results in Section V-B, we do not evaluate the No Variation Bridge policy, because we found it was unable to succeed at all with shifts for these factors in isolation, so there was no potential for composition.
While these policies worked for object/container position, we found they was unable to perform the task for the object orientation shifts we considered in isolation. We hypothesize this could be because different values for this factor have relatively small changes in their visual observations, but require significantly different behavior, which can make it challenging to learn when to apply the correct behavior. Therefore, for our evaluation on composing object orientation, we trained separate policies where we balance training batches such that 50% of in-domain data consists of data for object orientation.
We report these results in Table VI. We find that with prior data, object/container position achieves similar composition as our original object position factor, with a success rate of 20/36 for both. However, when training from scratch, the policy is able to compose this new factor more effectively than the original, achieving a success rate of 10/36 compared to 1/36. This could be because different values for object/container position are more visually distinguishable, and their required behavior is also significantly more different, which could make it easier to learn when to apply the correct behavior.
Our policies can sometimes compose object orientation, although composition for this is the weakest compared to the other factors we study. This could be due to the aforementioned challenges with learning for this factor, as well as because our data balancing may have insufficiently represented the other factors.
Figure 17: Per-factor value success rates of policies from our pairwise composition results in Table I. Factor values are placed in tiers, where increasing tiers are more dissimilar from the base factor values $f^{*}$. Composition is generally more challenging for factor values that are more dissimilar from the base factor values.
C-E Factor Similarity Analysis
Here, we provide additional analysis on when our policies are able to compose factor values from our pairwise evaluation in Section V-B. To do this, for each factor, we consider how similar each of the non-base factor values we consider are to the base factor value, with respect to a policy’s ability to generalize across factor values.
To compute a similarity metric, we consider the success rates from our results in Table I, in particular for the policy trained on No Variation data and BridgeData V2. We then obtain a success rate for each non-base factor value, by aggregating the results for all factor value pairs that contain that factor value. We use this success rate as our similarity metric, where higher success rates indicate greater similarity, because this captures how well a policy trained on base factor values generalizes to other factor values.
Factor Tier 1 Tier 2 Tier 3
Object Pos Down (3) Up (4) Left (2)
Object Type Wooden (3) Gray (4) Plastic (2)
Container Type Blue Plate (4) Pink Bowl (3) White Cup (2)
Table Height Higher 5cm (2) Lower 5cm (3) Lower 8cm (4)
Table Tex Brown Wood (2) Gift Wrap (4) White Marble (3)
TABLE VII: Tiers for each factor value used in Fig. 17, determined using our success rate-based similarity metric as described in Section C-E. In parentheses next to each factor, we provide the row number where the factor value is visualized in Fig. 7.
Using this similarity metric, we rank the non-base factor values for each factor as Tier 1, Tier 2, or Tier 3, where a higher tier is more dissimilar from the base factor value. We list the tiers for each factor value in Table VII. We then take the aggregated per-factor value success rates for each policy from our results in Table I (where we aggregate success rates as we do for computing the similarity metric), and plot this against our factor value tiers in Fig. 17.
We find that the policies trained on L data (orange and pink lines) generally achieve lower compositional success rates for factor values that are more dissimilar from the base factor values $f^{*}$, suggesting that composition is more challenging for these more dissimilar values, although this trend is not strictly monotonic. We note that part of this effect could be because our success rate-based similarity metric may also capture how challenging factor values are in general.
C-F R3M
@@
C-G VC-1
We train and evaluate VC-1 policies using the same procedure as with R3M, except using the ViT-L version of VC-1 instead of R3M.